### v1

@@ 3.5 Training
We train FLIP only on the Segment Anything Dataset, performing approximately 1.5 M updates at a batch size of 256, covering about 35.84 % of the dataset. Each training example has an associated object mask, from which we compute a 2D Gaussian distribution capturing the object’s center, scale, and orientation. To improve generalization, we apply random perturbations to this Gaussian by slightly shifting $x$ and $y$, stretching or compressing $\sigma_{x}$ and $\sigma_{y}$, and introducing small rotations. Tokens are sampled directly from this perturbed 2D Gaussian, with the total number of tokens, $N$, varied around a predefined mean $\mu$ during training to further improve generalization.
The training consists of three phases:
  • Phase 1: Training begins with a mean token count $\mu=128$ and half-resolution inputs for 800k steps.
  • Phase 2: The token count increases to $\mu=512$ with full-resolution inputs for the next 600k steps.
  • Phase 3: A final fine-tuning phase uses $\mu=4096$ tokens for 100k steps, matching the number of patches used by SAM.
We train three model sizes FLIP-S (2.6M parameters), FLIP-M (6.5M parameters), and FLIP-L (26.5M parameters). Notably, these models are significantly smaller in parameter count than SAM variants (up to 641.1M) and FastSAM-x (72.2M), with only FastSAM-s falling in a similar range (11.8M).
We optimize three losses to predict accurate masks, bounding boxes, and positional ($\mathbf{z}_{\text{pos}}$) codes.
  • Mask Loss: Binary cross-entropy loss between predicted mask logits $\hat{M}$ and ground truth $M$:
    $s_{\text{mask},i}=\frac{1}{\sum_{j=1}^{N_{i}}M_{ij}}$
    $\mathcal{L}_{\text{mask}}=\frac{1}{B}\sum_{i=1}^{B}s_{\text{mask},i}\sum_{j=1}^{N_{i}}\ell_{\text{BCE}}(\hat{M}_{ij},M_{ij}),$
    where $B$ is the batch size.
  • Position Loss: Penalizes deviations in mean position ($\mu$), scale ($\sigma$), and rotation ($\theta$):
    $\mathcal{L}_{\text{pos}}=\frac{1}{B}\sum_{i=1}^{B}\bigl{(}\ \frac{1}{\sigma_{i}}|\hat{\mu}_{i}-\mu_{i}\|^{2}+\|\hat{\sigma}_{i}-\sigma_{i}\|^{2}+\|\hat{\theta}_{i}-\theta_{i}\|^{2}\bigr{)}.$
  • Bounding Box Loss: Matches predicted bounding boxes $\hat{B}_{i}$ to ground truth $B_{i}$:
    $\mathcal{L}_{\text{bbox}}=\frac{1}{B}\sum_{i=1}^{B}\frac{1}{\sigma_{i}}\|\hat{B}_{i}-B_{i}\|^{2}.$
Total Loss
The overall loss combines mask, position, and bounding box losses:
$\mathcal{L}_{\text{stage1}}=\mathcal{L}_{\text{mask}}+\lambda_{\text{pos}}\mathcal{L}_{\text{pos}}+\lambda_{\text{bbox}}\mathcal{L}_{\text{bbox}},$
where $\lambda_{\text{pos}}$ and $\lambda_{\text{bbox}}$ control weightings.
4 Results
We evaluate FLIP on our newly constructed synthetic dataset ObjaScale, which we designed to stress-test scale invariance, and on three standard benchmarks: Hypersim, KITTI-360, and OpenImages Roberts et al. (2021); Liao et al. (2022); Kuznetsova et al. (2020). Notably, FLIP was neither trained nor fine-tuned on the ObjaScale dataset. Our experiments compare FLIP against two state-of-the-art segmentation methods, SAM Kirillov et al. (2023) and FastSAM Zhao et al. (2023), focusing on handling small objects, achieving high IoU, and balancing parameter/runtime trade-offs.
@@ 4 Results
(b) FastSAM-x
(c) FLIP-L
@@
Figure 7: IoU (ObjaScale) heatmaps illustrating relative vs. absolute mask size. FLIP-L retains strong accuracy even for small objects, provided there are enough pixels (e.g., $\geq 10\times 10$).
Table 1: Comparison of Mean IoU (%) and Std IoU (%) across different datasets.
@@
Model Size (M) Time (ms) Hypersim KITTI-360 OpenImage ObjaScale (ours)
Pre/Pos Model Total Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%)
FastSAM-s 11.8 4.56 4.84 9.40 36.80 ± 34.62 37.71 ± 32.76 61.14 ± 32.10 48.20 ± 35.69
FastSAM-x 72.2 4.54 18.46 23.00 43.39 ± 35.95 38.92 ± 34.06 69.31 ± 29.25 47.11 ± 36.41
FLIP-S (ours) 2.1 22.58 9.63 32.21 61.51 ± 27.16 59.30 ± 19.58 70.75 ± 23.51 77.69 ± 19.83
FLIP-M (ours) 6.5 23.84 15.04 38.88 64.36 ± 28.32 61.45 ± 19.64 76.76 ± 21.73 82.71 ± 18.24
FLIP-L (ours) 26.5 22.54 18.09 40.63 66.83 ± 27.74 61.85 ± 19.56 78.37 ± 21.77 84.94 ± 17.16
SAM-B 93.7 11.06 63.40 74.46 71.46 ± 20.88 62.38 ± 21.41 84.72 ± 15.38 71.38 ± 25.36
SAM-L 312.3 11.07 139.17 150.24 72.13 ± 21.21 62.73 ± 20.31 86.94 ± 13.41 72.68 ± 25.22
SAM-H 641.1 10.96 222.00 232.96 72.37 ± 21.65 62.47 ± 20.52 87.06 ± 13.53 73.76 ± 24.59
4.1 Experimental Setup and Dataset Creation
We selected 68 diverse categories from Objaverse Deitke et al. (2023) and combined each with high-resolution HDRI Haven Zaal, Greg backgrounds in Blender, yielding 10,200 synthetic images, which form our ObjaScale dataset. Object scale and image resolution (ranging from $512$ to $8192$) were randomized, producing masks spanning minuscule ($<\!0.0001\%$) to large fractions of the image (Fig. 8). For SAM and FastSAM, bounding-box prompts computed from ground-truth masks were used; FLIP employed 2D Gaussian prompts by design.
4.2 Performance vs. Relative Mask Size
On the ObjaScale dataset, Figure 10 plots IoU against relative mask size. FLIP-L consistently surpasses the other models for small objects ($<0.1\%$ of the image). FastSAM-x drops steeply, while SAM-H tapers more moderately. By adaptively sampling around the actual pixel area, FLIP-L retains strong IoU scores even when objects occupy a tiny fraction of the scene.
4.3 Heatmap Analysis and Small-Object Segmentation
Figure 7 highlights segmentation performance with respect to both relative and absolute mask size on ObjaScale. SAM-H struggles when the mask ratio diminishes, and FastSAM-x fails below a mask to image ratio of $0.01\%$. FLIP-L, by contrast, primarily depends on whether the object covers enough pixels to support a sufficiently detailed patch representation.
4.4 Evaluation on Benchmark Datasets
Table 1 compares Mean and Std IoU across Hypersim, KITTI-360, OpenImages and ObjaScale. FLIP exceeds the best FastSAM variant and remains close to SAM, while requiring substantially fewer parameters and competitive inference time. Notably, this is despite its input patch sampling and output patch selection (pre-/pos-processing times) are not yet optimized for inference time.
@@ 4.4 Evaluation on Benchmark Datasets
4.5 Varying the Patch Token Budget
Lastly, we evaluated how limiting FLIP’s patch sampling impacts performance. Figure 9 shows FLIP-L on OpenImages for different token counts; even with only 100 patches, IoU remains near its peak, underscoring robust efficiency under constrained computational budgets.
Figure 9: FLIP-L performance on OpenImages under different numbers of sampled patches.
Overall, these results highlight FLIP’s ability to accurately segment objects across scales, remain parameter-efficient, and adapt to variable computational budgets while matching or exceeding competing methods.
Figure 10: Intersection over Union (IoU) vs. relative mask size. FLIP maintains high segmentation accuracy even at extreme scales. Shaded regions show 95% confidence intervals.

### v2

@@ 3.4 Training
We train FLIP exclusively on META’s SA-1B Dataset Kirillov et al. (2023) with four model variants: FLIP-Tiny (0.51M parameters), FLIP-Small (2.3M parameters), FLIP-Middle (11.5M parameters), and FLIP-Large (96.6M parameters), performing approximately 8.5M, 7.4M, 5.8M and 2.7M updates respectively with a batch size of 256. These models are significantly more parameter-efficient than SAM variants (up to 641.1M parameters).
For data augmentation, we apply random perturbations to the 2D Gaussian derived from ground truth masks by shifting center coordinates, stretching or compressing $\sigma_{x}$ and $\sigma_{y}$, and introducing small rotations. The number of sampled tokens $N$ is varied around $\mu=512$ during training to improve generalization.
Sparse Mask Prediction.
Instead of dense mask prediction, we employ a computationally efficient sparse sampling strategy that focuses on boundary regions. Pixels are grouped into seven distance-based categories according to their L1 distance from mask boundaries: [0-2), [2-4), [4-8), [8-16), [16-32), [32-64), and [64$\infty$) pixels. For each sample, we select 2048 pixels distributed across these groups with equal sampling from inside and outside the mask to maintain class balance.
We implement adaptive sampling that dynamically redistributes pixels across distance groups based on prediction difficulty. Groups with lower IoU performance receive more samples in subsequent iterations, ensuring computational focus on challenging boundary regions.
The sparse mask loss reduces computational complexity from $O(HW)$ to $O(k)$ where $k\ll HW$:
$\mathcal{L}=\text{BCE}(\hat{M}_{\text{sparse}},M_{\text{sparse}})$ (1)
where $\hat{M}_{\text{sparse}}$ and $M_{\text{sparse}}$ represent predicted and ground truth values at sampled locations.
3.5 Inference
During inference, FLIP employs two strategies to efficiently generate segmentation masks while maintaining the sparse computation paradigm.
5-Sigma Bounding Box Prediction
Instead of predicting mask values for all pixels in the image, we exploit the Gaussian prior to restrict predictions to a 5-sigma bounding box around the object center. Given the Gaussian parameters $(\mu_{x},\mu_{y},\sigma_{x},\sigma_{y})$ from the input prompt, we compute a bounding box using $\sigma_{\text{iso}}=\max(\sigma_{x},\sigma_{y})$ and sample pixels only within the region $[\mu_{x}\pm 5\sigma_{\text{iso}},\mu_{y}\pm 5\sigma_{\text{iso}}]$. This reduces the number of mask queries from $O(HW)$ to $O(k^{2})$ where $k\propto\sigma_{\text{iso}}$, providing significant speedup for small objects while maintaining full accuracy for larger once.
Hierarchical Refinement
For further efficiency, we introduce a hierarchical inference strategy that progressively refines the mask prediction. Starting from a coarse resolution, we:
  1. 1.
    Predict initial mask values on the coarse grid
  2. 2.
    Identify uncertain regions where predictions fall inside $[\tau,1-\tau]$ (default $\tau=0.01$)
  3. 3.
    Upsample the mask by a factor $\alpha$ (default $\alpha=4$) using bilinear interpolation
  4. 4.
    Re-query only the uncertain pixels at the higher resolution
  5. 5.
    Repeat until reaching the target resolution
This approach dramatically reduces computational cost by focusing refinement on ambiguous boundary regions. The number of refinement steps $n$ is automatically determined as $n=\lfloor\log_{\alpha}(k_{\text{target}}/k_{\text{init}})\rfloor$, where $k_{\text{init}}$ and $k_{\text{target}}$ are the initial and target resolutions.
Both strategies can be combined, first using the 5-sigma bounding box to limit the spatial extent, then applying hierarchical refinement within that region. This enables real-time mask generation even for high-resolution images while maintaining maximum segmentation quality.
@@ 3.5 Inference
(b) EfficientSAM-S
(c) MobileSAM
(d) FastSAM-x
(e) FLIP-Large
@@
Figure 7: IoU (ObjaScale) heatmaps illustrating relative vs. absolute mask size. FLIP-Large retains strong accuracy even for small objects, provided there are enough pixels (e.g., $\geq 10\times 10$).
Figure 8: Comparison of mean IoU across six diverse datasets (Hypersim, KITTI-360, OpenImages, COCO, LVIS, ObjaScale) plotted against inference time (in milliseconds). FLIP variants consistently outperform SAM variants while being significantly faster.
Table 1: Comparison of Mean IoU (%) and Std IoU (%) across different datasets
@@
Model Size (M) Time (ms) Hypersim KITTI-360 OpenImages COCO LVIS ObjaScale (ours) Average
Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%)
FastSAM-s 11.8 9.94 36.80 ± 34.62 37.71 ± 32.76 61.14 ± 32.10 46.51 ± 33.73 37.11 ± 35.19 48.20 ± 35.69 44.58 ± 34.02
FastSAM-x 72.2 24.32 34.36 ± 34.90 38.92 ± 34.06 69.31 ± 29.25 55.31 ± 34.55 43.24 ± 37.47 47.11 ± 36.41 48.04 ± 34.44
MobileSAM 10.13 21.15 68.03 ± 21.36 64.64 ± 18.87 82.20 ± 16.74 73.80 ± 17.48 71.59 ± 21.36 67.72 ± 25.03 71.33 ± 20.14
EfficientSAM-T 10.22 26.75 68.35 ± 24.52 60.82 ± 24.72 84.32 ± 15.08 75.93 ± 17.04 75.56 ± 19.95 68.74 ± 26.43 72.29 ± 21.29
EfficientSAM-S 26.41 47.98 69.65 ± 22.20 64.88 ± 19.35 85.99 ± 14.54 77.07 ± 16.81 75.18 ± 21.06 67.80 ± 26.76 73.43 ± 20.12
SAM-B 93.7 72.67 71.46 ± 20.88 62.38 ± 21.41 84.72 ± 15.38 76.07 ± 16.95 76.93 ± 19.17 71.38 ± 25.36 73.82 ± 19.86
SAM-L 312.3 148.78 72.13 ± 21.21 62.73 ± 20.31 86.94 ± 13.41 78.19 ± 16.05 77.93 ± 19.43 72.68 ± 25.22 75.10 ± 19.27
SAM-H 641.1 232.04 72.37 ± 21.65 62.47 ± 20.52 87.06 ± 13.53 78.41 ± 16.15 78.36 ± 19.50 73.76 ± 24.59 75.41 ± 19.32
FLIP-Tiny 0.51 9.82 73.17 ± 19.36 70.04 ± 15.27 83.83 ± 12.72 77.76 ± 13.55 76.88 ± 18.33 87.76 ± 14.93 78.24 ± 15.69
FLIP-Small 2.3 12.19 73.07 ± 19.80 69.91 ± 15.69 86.65 ± 11.27 79.23 ± 13.32 77.92 ± 18.34 88.96 ± 15.14 79.29 ± 15.59
FLIP-Middle 11.5 17.54 73.68 ± 20.13 69.86 ± 15.76 87.89 ± 10.08 80.16 ± 13.15 78.88 ± 18.31 89.08 ± 15.04 79.93 ± 15.41
FLIP-Large 96.6 38.65 73.66 ± 20.51 69.75 ± 16.03 89.33 ±   8.76 80.79 ± 13.06 79.42 ± 18.13 89.02 ± 15.21 80.33 ± 15.28
FLIPh-Tiny 0.51 9.05 71.36 ± 18.65 68.53 ± 15.08 83.03 ± 12.65 76.57 ± 13.25 75.41 ± 17.37 85.72 ± 14.52 76.77 ± 15.25
FLIPh-Small 2.3 10.94 71.93 ± 18.99 68.34 ± 15.57 85.73 ± 11.29 78.15 ± 13.04 76.71 ± 17.41 86.97 ± 14.68 77.97 ± 15.16
FLIPh-Middle 11.5 14.34 72.65 ± 19.21 68.28 ± 15.68 86.98 ± 10.01 78.98 ± 12.89 77.58 ± 17.25 86.95 ± 14.68 78.57 ± 14.95
FLIPh-Large 96.6 29.36 73.19 ± 19.56 68.05 ± 16.18 88.28 ±   8.96 79.67 ± 12.83 78.40 ± 17.10 87.36 ± 14.61 79.16 ± 14.87
Table 2: FLIP-Small ablation study
Ablation Params (M) Time (ms) Hypersim KITTI-360 OpenImages COCO LVIS ObjaScale (ours) Average
Mean ± Std Mean ± Std Mean ± Std Mean ± Std Mean ± Std Mean ± Std Mean ± Std
Baseline 2.3 12.19 73.07 ± 19.80 69.91 ± 15.69 86.65 ± 11.27 79.23 ± 13.32 77.92 ± 18.34 88.96 ± 15.14 79.29 ± 15.59
Baselineh 2.3 10.94 71.93 ± 18.99 68.34 ± 15.57 85.73 ± 11.29 78.15 ± 13.04 76.71 ± 17.41 86.97 ± 14.68 77.97 ± 15.16
InitialPosEmb 2.2 10.37 73.41 ± 19.73 68.96 ± 15.48 86.29 ± 11.54 79.22 ± 13.14 78.42 ± 17.98 89.05 ± 14.98 79.23 ± 15.47
InitialPosEmbh 2.2 8.98 72.03 ± 18.92 67.36 ± 15.27 85.45 ± 11.51 77.94 ± 12.98 76.92 ± 17.15 86.84 ± 14.63 77.76 ± 15.08
LaPE-SNorm 2.3 10.88 73.86 ± 19.37 69.92 ± 15.20 86.41 ± 11.40 79.29 ± 13.12 78.42 ± 18.00 89.12 ± 14.77 79.50 ± 15.31
LaPE-SNormh 2.3 8.79 71.03 ± 17.74 65.87 ± 13.78 80.20 ± 11.98 74.42 ± 12.71 74.34 ± 16.30 83.01 ± 14.61 74.81 ± 14.52
MaxPatch4x4 1.7 11.46 72.69 ± 19.71 69.43 ± 15.57 86.01 ± 11.55 78.73 ± 13.28 77.55 ± 18.30 88.48 ± 14.96 78.81 ± 15.56
MaxPatch4x4h 1.7 10.09 71.45 ± 18.97 67.87 ± 15.46 85.11 ± 11.47 77.61 ± 13.07 76.31 ± 17.43 86.55 ± 14.57 77.48 ± 15.16
MaxPatch64x64 3.0 12.54 73.45 ± 19.52 69.78 ± 15.93 86.67 ± 11.37 79.33 ± 13.15 78.39 ± 18.02 89.04 ± 14.89 79.44 ± 15.48
MaxPatch64x64h 3.0 11.10 72.19 ± 18.80 68.21 ± 15.80 85.73 ± 11.32 78.16 ± 12.91 77.04 ± 17.08 87.05 ± 14.41 78.06 ± 15.05
MinPatch4x4 2.2 11.45 69.39 ± 23.77 70.30 ± 15.60 86.42 ± 11.60 77.27 ± 16.62 72.08 ± 25.85 88.14 ± 17.28 77.27 ± 18.45
MinPatch4x4h 2.2 10.07 68.53 ± 22.58 68.82 ± 15.46 85.53 ± 11.60 76.31 ± 16.23 71.07 ± 24.74 86.05 ± 16.78 76.05 ± 17.90
NoDynamicSampling 2.3 12.17 71.03 ± 20.14 70.03 ± 15.46 87.47 ± 11.07 79.04 ± 13.65 76.40 ± 19.00 87.82 ± 15.95 78.63 ± 15.88
NoDynamicSamplingh 2.3 10.69 70.57 ± 19.46 68.54 ± 15.26 86.63 ± 11.07 78.29 ± 13.38 75.89 ± 18.11 86.66 ± 15.20 77.76 ± 15.41
Table 3: Segmentation Accuracy on small-scale objects ($<1$% image area)
Model Size (M) Time (ms) Hypersim KITTI-360 OpenImages COCO LVIS ObjaScale (ours) Average
Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%) Mean ± Std IoU (%)
FastSAM-s 11.8 10.02 31.91 ± 32.85 32.55 ± 31.32 50.59 ± 32.66 38.50 ± 32.74 31.12 ± 33.01 45.91 ± 35.16 38.43 ± 32.96
FastSAM-x 72.2 24.37 24.77 ± 30.20 33.12 ± 33.34 56.58 ± 31.82 44.09 ± 34.05 34.96 ± 35.00 44.13 ± 35.62 39.61 ± 33.34
MobileSAM 10.13 21.83 65.67 ± 20.77 62.36 ± 18.41 76.76 ± 19.01 70.19 ± 17.36 68.46 ± 21.25 65.91 ± 24.86 68.22 ± 20.28
EfficientSAM-T 10.22 26.75 68.61 ± 20.18 63.11 ± 17.56 79.67 ± 16.78 72.91 ± 16.41 72.79 ± 19.80 67.11 ± 26.17 70.70 ± 19.48
EfficientSAM-S 26.41 47.98 66.74 ± 21.33 63.64 ± 18.01 80.82 ± 17.48 73.26 ± 16.78 71.64 ± 21.16 65.86 ± 26.63 70.33 ± 20.23
SAM-B 93.7 73.33 69.97 ± 19.58 63.05 ± 18.28 81.08 ± 16.87 74.48 ± 15.60 74.53 ± 19.11 69.79 ± 25.41 72.15 ± 19.14
SAM-L 312.3 149.45 69.76 ± 20.29 62.40 ± 18.44 82.76 ± 15.33 75.39 ± 15.59 74.75 ± 19.71 71.12 ± 25.32 72.70 ± 19.11
SAM-H 641.1 232.67 70.12 ± 20.66 62.02 ± 18.50 82.66 ± 15.81 75.62 ± 15.63 75.18 ± 19.83 72.26 ± 24.71 72.98 ± 19.19
FLIP-Tiny 0.51 8.58 71.04 ± 19.97 67.92 ± 15.59 81.53 ± 14.35 76.57 ± 13.93 75.18 ± 18.86 87.53 ± 15.16 76.63 ± 16.31
FLIP-Small 2.3 10.51 69.85 ± 20.17 67.53 ± 15.55 84.13 ± 12.70 77.18 ± 13.76 75.46 ± 18.89 88.75 ± 15.29 77.15 ± 16.06
FLIP-Middle 11.5 13.93 69.77 ± 20.39 66.69 ± 15.32 85.40 ± 11.21 77.69 ± 13.67 75.97 ± 18.98 88.86 ± 15.22 77.39 ± 15.80
FLIP-Large 96.6 29.89 69.18 ± 20.58 66.01 ± 15.39 86.85 ±   9.31 77.83 ± 13.47 76.12 ± 18.72 88.67 ± 15.51 77.44 ± 15.50
FLIPh-Tiny 0.51 8.21 68.94 ± 19.00 66.07 ± 15.14 80.26 ± 14.07 74.72 ± 13.31 73.45 ± 17.57 85.45 ± 14.72 74.81 ± 15.63
FLIPh-Small 2.3 10.01 68.71 ± 19.19 65.64 ± 15.17 82.58 ± 12.71 75.54 ± 13.10 74.09 ± 17.66 86.72 ± 14.79 75.55 ± 15.43
FLIPh-Middle 11.5 12.82 68.83 ± 19.34 64.78 ± 14.94 83.82 ± 11.08 75.92 ± 13.01 74.54 ± 17.55 86.71 ± 14.81 75.77 ± 15.12
FLIPh-Large 96.6 26.81 68.99 ± 19.63 63.94 ± 15.19 85.11 ±   9.47 76.21 ± 12.75 75.06 ± 17.40 86.99 ± 14.86 76.05 ± 14.88
@@
4 Results
We evaluate FLIP on five standard benchmarks: Hypersim, KITTI-360, OpenImages, COCO and LVIS Roberts et al. (2021); Liao et al. (2022); Kuznetsova et al. (2020); Lin et al. (2014); Gupta et al. (2019). Additionally, we evaluate FLIP on our newly constructed synthetic dataset ObjaScale, designed to stress-test scale invariance. Like SAM, FLIP was trained on the SA-1B dataset Kirillov et al. (2023), but was neither trained nor fine-tuned on the ObjaScale or any of the other evaluation datasets. Our experiments compare FLIP against state-of-the-art segmentation methods namely SAM Kirillov et al. (2023), EfficientSAM Xiong et al. (2023), MobileSAM Zhang et al. (2023), and FastSAM Zhao et al. (2023). Besides comparing the overall performance, we focus on handling small objects, achieving high IoU, and balancing parameter/runtime trade-offs.
4.1 Experimental Setup and Dataset Creation
We selected 68 diverse categories from Objaverse Deitke et al. (2023) and combined each with high-resolution HDRI Haven Zaal, Greg backgrounds in Blender, yielding 10,200 synthetic images, which form our ObjaScale dataset. Image resolution was randomized from $512$ to $8192$ pixels, while object dimensions were constrained to a maximum width/height of $256$ pixels, producing masks spanning minuscule ($<\!0.0001\%$) to large fractions of the image area (up to $25\%$) (Figure 9). For SAM variants, bounding-box prompts computed from ground-truth masks were used; FLIP employed 2D Gaussian prompts by design.
4.2 Overall Segmentation Performance
We denote hierarchical inference variants with the subscript “h” (e.g. FLIPh-Small); they share weights with their standard counterparts but perform progressive mask refinement (see subsection 3.5).
Table 1 presents comprehensive results across all datasets. FLIP consistently achieves superior performance compared to SAM variants while using substantially fewer parameters. FLIP-Large attains 80.33% mean IoU with 96.6M parameters, outperforming SAM-H’s 75.41% IoU that requires 641.1M parameters—a 6.6$\times$ parameter reduction with improved accuracy. Even our smallest model, FLIP-Tiny with only 0.51M parameters, achieves 78.24% mean IoU, surpassing all SAM variants including the 641M parameter SAM-H.
The performance advantage is particularly pronounced on our ObjaScale dataset, where FLIP models achieve 87–89% IoU compared to SAM-H’s 73.76%. This demonstrates FLIP’s superior scale invariance and ability to handle objects across diverse size ranges. On established benchmarks, FLIP maintains competitive performance: achieving 89.33% IoU on OpenImages (vs. SAM-H’s 87.06%) and 80.79% on COCO (vs. SAM-H’s 78.41%).
4.3 Parameter Efficiency and Performance
Figure 1 illustrates the parameter efficiency of FLIP variants. While maintaining superior segmentation quality, FLIP models operate with orders of magnitude fewer parameters than SAM variants. The efficiency gains translate to faster inference times (Figure 8), with FLIP-Large requiring only 38.65ms compared to SAM-H’s 232.04ms, representing a 6$\times$ speedup alongside improved accuracy (measured on a Nvidia RTX4090 graphics card).
Hierarchical Inference Trade-offs.
The hierarchical inference variants (FLIPh) demonstrate the effectiveness of our progressive refinement strategy, achieving comparable performance with reduced computational cost. Comparing standard inference with hierarchical variants reveals a general pattern: hierarchical inference achieves 98–99% of the original performance while providing $\approx$ 8–24% speed improvements. For instance, FLIPh-Large maintains 79.16% mean IoU while reducing inference time to 29.36ms. This trade-off is particularly favorable for real-time applications where slight accuracy reductions are acceptable for substantial computational savings.
4.4 Small Object Segmentation
Table 3 reveals FLIP’s capability for small object segmentation (objects $<$1% image area). FLIP-Large achieves 77.44% mean IoU on small objects compared to SAM-H’s 72.98%, demonstrating the effectiveness of our fovea-like patching mechanism. This advantage is most pronounced on ObjaScale, where FLIP-Tiny achieves 87.53% IoU on small objects compared to SAM-H’s 72.26%.
The heat-map analysis (Figure 7) offers clear evidence of FLIP’s scale-invariant behavior: whereas SAM variants—and FastSAM-x in particular—suffer sharp performance drops when the mask-to-image ratio falls below 0.01%, FLIP’s IoU remains stable until the object’s absolute footprint shrinks to roughly $10\times 10$ pixels. Because FLIP’s accuracy is governed by absolute pixel coverage rather than relative image area, it robustly segments small objects even in ultra-high-resolution scenes—a critical advantage for real-world applications.
4.5 Architectural Robustness and Ablation Analysis
Our ablation study (Table 2) reveals that FLIP is remarkably robust to architectural variations while certain design choices provide measurable improvements. The baseline FLIP-Small (79.29% mean IoU) serves as our reference point for evaluating architectural components.
Positional Embedding Strategies.
The comparison between baseline and “InitialPosEmb” (79.23% vs. 79.29%) shows that our layer-wise positional embedding injection provides marginal but consistent benefits over initial embedding only. However, the “LaPE-SNorm” variant, which applies shared normalization across positional embeddings, achieves the best performance at 79.50% mean IoU, while underperforming in the hierarchical inference setting.
Patch Resolution Configurations.
Testing different patch size ranges reveals interesting trade-offs. Restricting maximum patch size to 4$\times$4 (“MaxPatch4x4”) reduces parameters to 1.7M with only modest performance degradation (78.81% IoU), demonstrating FLIP’s efficiency even with limited resolution diversity. Conversely, extending to 64$\times$64 patches (“MaxPatch64x64”) slightly improves performance (79.44%) at the cost of increased parameters (3.0M), suggesting diminishing returns from very large patches.
Removing the smallest 1$\times$1 patches (“MinPatch4x4”) significantly impacts performance on LVIS (72.08% vs. 77.92%), highlighting the importance of fine-grained detail capture for complex segmentation tasks while maintaining strong performance on other datasets.
Sampling Strategy Impact.
The “NoDynamicSampling” ablation (78.63% vs. 79.29%) demonstrates that our adaptive pixel sampling strategy provides consistent benefits across datasets. Without dynamic sampling, each distance group uses the same number of pixels rather than adaptively redistributing based on prediction difficulty. The performance gap is particularly notable on Hypersim (71.03% vs. 73.07%), where focusing computational resources on challenging boundary regions better handles the diverse object shapes in simulated indoor scenes.
4.6 Cross-Dataset Generalization
FLIP demonstrates robust generalization across diverse evaluation benchmarks without dataset-specific fine-tuning. The performance gains are consistent across synthetic (Hypersim), automotive (KITTI-360), and large-scale natural image datasets (OpenImages, COCO, LVIS). Notably, FLIP-Tiny achieves 70.04% IoU on KITTI-360 compared to SAM-H’s 62.47%, despite using 1,257$\times$ fewer parameters. This 7.57 percentage point improvement highlights how FLIP’s fovea-like attention mechanism effectively handles automotive scenarios characterized by numerous small, distant objects.
The architectural ablations confirm that FLIP’s performance is remarkably robust to design variations. While removing the smallest patches (MinPatch4x4) impacts performance on complex datasets like LVIS, and the adaptive pixel sampling during training (NoDynamicSampling ablation) provides consistent improvements, the overall architecture maintains strong performance across different configurations. This robustness suggests that FLIP’s core fovea-like attention mechanism, rather than any single architectural choice, drives its superior accuracy, particularly for challenging small objects.